OpenStudy (anonymous):

if 1/(a+b) + 1/(c+b)=1/b........then prove a, b, c are in geometric progression

OpenStudy (anonymous):

solve the equation by bringing all terms in numerator .. you would get ac + ab + bc+ b^2 = bc + b^2 + ab + b^2 simplify to get ac = b^2 hence proved

OpenStudy (anonymous):

thanx mate'

OpenStudy (anonymous):

a geometric progression, also known as a geometric sequence, is a sequence of numbers where each term after the first is found by multiplying the previous one by a fixed non-zero number called the common ratio. For satisfying G.P., three consecutive terms a ,b and c shall satisfy the following equation: b^2 = ac